OpenStudy (anonymous):

verify the identities: 1. x=60 degrees, sin 2x=2 sin x cos x 2. cos (2theta)/sin theta cos theta = cot theta minus tan theta

OpenStudy (anonymous):

in (1), just substitute x=60 deg to the given identity of sin 2x...

OpenStudy (anonymous):

sin 2x = 2 sin x cos x is already an identity derived as double angle formula... using sin (x + x) = sin x cos x + sin x cos x = 2 sin x cos x

OpenStudy (anonymous):

in (2) we use double angle formula for cos 2theta... \[\cos 2\theta=\cos^2\theta-\sin^2\theta\]

OpenStudy (anonymous):

\[\frac{ \cos2\theta }{ \sin \theta \cos \theta }=\cot \theta-\tan \theta\] \[\frac{ \cos^2\theta }{ \sin \theta \cos \theta }-\frac{ \sin^2\theta }{ \sin \theta \cos \theta }=\cot \theta-\tan \theta\] \[\frac{ \cos \theta }{ \sin \theta }-\frac{ \sin \theta }{ \cos \theta }=\cot \theta-\tan \theta\] that proves the identity for (2)...

OpenStudy (anonymous):

since \[\cot \theta = \frac{ \cos \theta }{ \sin \theta }\] and \[\tan \theta = \frac{ \sin \theta }{ \cos \theta }\]